OpenStudy (anonymous):

Solve Z = 9x + 9y for y. y = P – 9x y = Z – 9x – 9 y= z-x/9 y= z -9x/9

OpenStudy (anonymous):

@Valkarie70

OpenStudy (anonymous):

ok, so this will be the last one i can answer for a little while because i got school todo also, but if you need help feel free to mention me and ill get right to it :) for this one you need to get y by itself to solf for it so lets start with that. Z=9x+9y Z-9x=9y (Z/9)-x=y because you divide both sides by 9 Answer: D) Y=(z-9x)/2 assuming this is supposed to have parenthesis.